# Bridging the Divide: Unraveling the Knowledge Gap in Data Visualization Research and Practice

### v2

@@ Abstract
Empirical research on perception and cognition has laid the foundation for visualization design, often yielding useful design guidelines for practitioners. However, it remains uncertain how well practitioners stay informed about such crucial visualization design knowledge. In this paper, we employed a mixed-method approach to explore the knowledge gap between visualization research and real-world design guidelines. We initially collected existing design guidelines from various sources and empirical studies from diverse publishing venues, analyzing their alignment and uncovering missing links and inconsistent knowledge. Subsequently, we conducted surveys and interviews with practitioners and researchers to gain further insights into their experiences and attitudes towards design guidelines and empirical studies, and their views on the knowledge gap between research and practice. Our findings highlight the similarities and differences in their perspectives and propose strategies to bridge the divide in visualization design knowledge.
Index Terms:
Data visualization, empirical research, design guideline, knowledge gap, survey, interview, content analysis.
@@ I Introduction
Data visualization has seamlessly integrated into our daily routines, serving as a crucial medium to convey intricate data across disciplines. Individuals encounter an array of data visualizations, spanning topics from climate change and election outcomes to pandemic health risks. Nowadays, people have a plethora of tools to create data visualizations, such as Tableau and PowerBI. Data visualization producers include not only scientists and analysts, but also journalists, managers, and even casual users. Despite their efforts to address various intricate design factors, such as perceptual effectiveness, aesthetics, and engagement, we still frequently come across bad or weird visualizations that result in heated discussions on social media [1].
Empirical research has provided foundational knowledge for data visualization. Graphical perception studies, beginning with the seminal work by Cleveland and McGill [2], have provided useful insights into which visual channels are more perceptually effective. Visual cognition studies have contributed to a better understanding of the higher-level processing of visualizations, including comprehension [3], engagement [4], memorability [5], and biases [6]. Guidelines for visualization design, such as including a zero baseline for a bar chart and using a bar chart instead of a pie chart for value comparison, are based on past empirical studies and are now frequently shared among practitioners [7]. Although empirical research continues to expand design knowledge and contextualize existing guidelines, the accessibility and application of this knowledge among visualization practitioners remain less understood.
In this study, we investigate the knowledge gap between visualization research and practice, aiming to understand their interplay better and promote evidence-based design practices. As the field of visualization expands and influences decision-making across society, understanding this gap is crucial. Our approach mirrors initiatives in fields like medicine, education, and policy making, where the shift has been from relying on tradition, intuition, and anecdotal experience to making decisions based on solid research findings [8, 9, 10, 11]. Our specific research questions include:
@@
  • What are the experiences and perceptions of both practitioners and researchers regarding the knowledge gap in research and practice?
@@
We employed a two-phase mixed-methods approach to explore these questions. We first collected data visualization design guidelines in the wild by conducting a keyword-based web search, and then conducted a qualitative content analysis to characterize their attributes. Our findings revealed a concentration of guidelines on basic charts and task types and identified three distinctive categories of the guidelines: problem-solving, enhancement, and alternative suggestion. To understand the agreement and discrepancies between guidelines and empirical research, we analyzed how they are aligned with each other; we focused on empirical studies that evaluate user performance across various visualization designs [12]. The alignment analysis revealed some inconsistencies between guidelines and their associated empirical evidence. Although mixed findings are typical in empirical research, this result implies that these nuances might not be communicated to practitioners. Furthermore, empirical studies not yet mapped to guidelines suggest new areas for guideline development, while guidelines without empirical backing offer avenues for fresh empirical research.
@@
The content analysis of the guideline landscape provided insights into the available design knowledge in practice, but it did not capture the lived experiences and perspectives of researchers and practitioners regarding the knowledge gaps. To gain a more comprehensive understanding, we supplemented this with surveys and interviews. Practitioners were mainly recruited from the Data Visualization Society [13], while researchers were recruited through community mailing lists. A total of 33 practitioners and 36 researchers filled out our survey questionnaire on guidelines, empirical studies, and knowledge transfer between the two communities. The survey results revealed similarities and differences in their access to guidelines, perceptions of their usefulness, and the challenges they encountered, also indicating a unified view on the lack of knowledge transfer between the two communities [14]. In follow-up interviews with ten selected participants from each community, we gained deeper insights into their experiences and struggles with guidelines, as well as their attitudes and perceived causes of the knowledge gap and ideas for addressing it.
@@
Our primary contribution lies in enhancing the understanding of the design knowledge gap between research and practice, paving the way for evidence-informed design decision-making. We synthesized our findings into a range of strategies and considerations, detailed in the discussion section, aimed at bridging the knowledge divide. These strategies include establishing consistent vocabularies for data visualizations, creating a triadic model that integrates guidelines, empirical evidence, and contextual examples, and improving guideline accessibility through detailed metadata annotation and template development. We also highlight the importance of fostering collective efforts in knowledge brokerage [15], essential for the growth and evolution of the field.
@@ II Background: Research, Practice, & Education
The field of data visualization has always been closely tied to practical needs and applications. Widely used software tools and programming toolkits, such as Tableau [16, 17], D3 [18], and Vega-Lite [19], began as research projects, as did visualization techniques such as treemaps [20] and parallel coordinates [21]. Researchers have also devoted much effort to understanding the challenges and struggles faced by practitioners in deriving new research insights [22, 23] and building new visualization tools for practitioners [24, 25]. Furthermore, the growing sub-fields of visual analytics [26] and design studies [27] directly address problems faced by domain experts across disciplines.
@@
Similarly, the visualization research community has developed theories and frameworks over the past few decades that provide established vocabularies for reading and creating visualizations, such as visual encoding channels, data types, and analytical tasks [28, 29]. To evaluate the usefulness of a visualization design, researchers conduct experiments that compare design alternatives. The results of these experiments provide guidance for selecting effective visual encoding channels [2, 30] and determining their effectiveness for different tasks [31, 32]. Some of these results are integrated into existing tools that automate visual encoding design [16, 33, 34]. Although these automated tools suffice for exploratory analysis, designers must still consider a range of design factors to ensure the successful and engaging communication of data, such as selecting appropriate color schemes and determining axis ranges.
@@
Researchers have developed pedagogical approaches to teaching data visualizations, such as the Five Design-Sheet Methodology [35] and VisItCards [36]. Other studies have examined data visualization literacy, which is defined as the ability to read information encoded in data visualizations and create them to answer questions about the data [37, 38, 39, 40, 41]. However, these works in pedagogy and literacy often overlook practical knowledge such as design guidelines, which are crucial for detailed design factors. For instance, pedagogical studies focus on teaching higher-level design processes and basic visualization vocabularies, whereas literacy studies center on evaluating chart reading tasks.
@@
Several recent studies have attempted to address more practical design needs. Diehl et al. launched VisGuides, a forum that provides practitioners with a central place to ask questions about the validity and quality of their visualizations and discuss guidelines around visualization design [42]. Wang et al. proposed a cheatsheet structure to introduce a visualization technique, including its construction process, similar techniques, and potential issues [43]. Likewise, other researchers have discussed guideline models for connecting with empirical studies [44] and describing their relationships across different stages in the design process [45].
@@
Similar efforts have also been made by practitioners as well. Resources such as From Data to Viz and DataVizProject provide user interfaces to select charts based on different data types and tasks [46, 47]. Companies publish design systems that often include recommendations on visualization design [48, 49], although they mostly focus on styling guidelines to match their brands.
@@
Currently, visualization design guidelines are dispersed across different resources, often through personal blogs and published books. It remains unclear how, or even if, practitioners learn critical design knowledge from these guidelines. Furthermore, researchers often convey research outcomes through blogs [50], podcasts [51], or social media in an ad-hoc manner. While two recent studies have begun investigating the longstanding gap between research and practice, they mostly focus on understanding holistic design processes rather than acquiring fundamental data visualization design knowledge [14, 52].
This paper aims to understand the potential knowledge gap between visualization design practice and research. We employ a range of mixed methods, including content analysis, surveys, and interviews, involving both practitioner and researcher groups. Our research shares a philosophy with fields such as education [53], medicine [11], and policy-making [9], which emphasize evidence-informed decision-making.
@@ III Guideline and Empirical Study Alignment
As an initial step toward understanding the knowledge gap, we investigated the range of visualization design guidelines available and how they align with empirical research in visualization, encompassing both contemporary and foundational studies.
@@ III-A1 Collecting guidelines
To ensure comprehensiveness in our collection, we extensively consulted a variety of guideline resources in practical settings, such as industry resources and online platforms. We first used web search by using a combination of different keywords, such as data visualization, guidelines, and similar synonyms, and went through up to five pages of the top results. We also looked at known resources such as blogs (e.g., DataWrapper [54]) and websites (e.g., from Data to Viz [46], Data Journalism [55]). Finally, we consulted the guideline collection curated by the Data Visualization Society [13], which provided additional blogs, websites, and books (e.g., How Charts Lie by Cairo [56]). Our search concluded once we reached a saturation of overlapping and duplicate guidelines, totaling 1,279 guidelines.
Next, we applied inclusion and exclusion criteria to serve our purpose of analyzing visualization design knowledge. For instance, we only considered guidelines relevant to data visualization design, adopting the definition from Chen et al. [57]; i.e., a guideline describes a process or a set of actions that may lead to a desired outcome or, alternatively, actions to be avoided to prevent an undesired outcome. Consequently, we excluded guidelines offering narrow styling advice, such as increasing font size. We also disregarded broader recommendations, such as making visualizations interactive. Similarly, we excluded generic chart suggestions, such as using a bar chart for a single series data, that are too evident to be considered guidelines. We also consolidated guidelines that significantly overlapped in content or purpose. After the filtering and aggregation process, we obtained the final collection of 390 guidelines from 32 diverse sources. Of these, 49% were derived from blog articles, 39% from style guidelines, and the remaining 12% from books.
@@
Our data collection focused specifically on guidelines readily accessible to practitioners, under the assumption, confirmed in Section IV), that they are unlikely to consult academic research papers. Therefore, we did not include design considerations from empirical papers, despite their relevance.
@@ III-A2 Collecting empirical studies
We used a combination of automated and manual searches to collect empirical papers. First, we initially used a keyword search (e.g., visualization, evaluation, study) on Scopus to collect papers from ACM CHI and EuroVis over the past decade. We collected the same data for IEEE VIS from VisPub [58]. Expanding on this initial collection, we supplemented it with recent papers published in ACM CHI, IEEE VIS, and Eurovis from 2021 to 2022 to ensure the dataset reflects the latest research. We also manually included relevant past papers that were not captured by our initial search from additional sources, including VisPerception [59] and a recent survey of perception-based visualization studies [60]; this process aided in gathering influential works published prior to 2011, such as the notable graphical perception experiment conducted by Cleveland & McGill [2]. Finally, two researchers cleaned up the data collection by removing any papers that did not present original empirical research and by only including studies that evaluate user performance on different visualization designs based on the characterization in a prior work [12] i.e., we excluded algorithm performance and system usability studies. The final collection of empirical studies consists of 235 research papers, spanning from 1911 to 2022. A detailed breakdown of the papers by venues and years is available in the supplementary material.
@@ III-B Analyzing Process
To facilitate content analysis [61] of the guidelines, two junior researchers compiled short descriptions from guideline articles into a spreadsheet. A lead researcher initially reviewed these descriptions and, when necessary, consulted the source articles. This process was to establish initial codes and code values, including chart types, data types, task types, and chart elements. The researcher drew upon established frameworks and vocabularies within the visualization community for guidance [29, 28, 46, 62, 63]. The two junior researchers then independently coded the data using the initial codebook. They expanded it as necessary, adding new codes related to purpose, problem, approach, and specific features like color, animation, 3D, and interaction, consulting with the lead researcher to resolve any questions or conflicts in coding. Lastly, the lead researcher thoroughly reviewed the final codes.
@@
We derived the one-sentence purpose of each study out of its abstract using GPT4 to facilitate the following alignment process. We initially attempted to codify aspects such as chart types and experiment methods. However, given the comprehensive empirical study landscape already explored in recent studies [64, 65, 60], we ultimately decided not to perform a detailed codification of the empirical studies. This approach allowed us to maintain a high-level analysis that aligns with our study’s strategic objectives, ensuring a focused and manageable research endeavor. Nonetheless, exploring finer-grained, code-level alignment in future work beyond our current entity-level alignment would be beneficial.

### v3

# Understanding the Research-Practice Gap in Visualization Design Guidelines

@@ Abstract
Although empirical research often underpins practical visualization guidelines, it remains unclear how well these research-driven insights are reflected in the guidelines practitioners actually use. In this paper, we investigate the research-practice gap in visualization design guidelines through a mixed-methods approach. We collected 390 design guidelines from practitioner-facing sources and 235 empirical studies to quantitatively assess their alignment. To complement this analysis, we conducted surveys with 69 participants (33 practitioners, 36 researchers) and in-depth interviews with 20 experts to examine their experiences, perceptions, and challenges. Our findings reveal discrepancies: empirical evidence often contradicts or only partially supports widely used guidelines, and the two communities prioritize different attributes of design. Based on these insights, we derive a holistic guideline template (integrating Context, Approach, Problem, and Purpose) and discuss actionable strategies, such as a triadic knowledge model.
@@ I Introduction
Empirical research has provided foundational knowledge for data visualization. Graphical perception studies, beginning with the seminal work by Cleveland and McGill [cleveland1984graphical], have provided useful insights into which visual channels are more perceptually effective. Visual cognition studies have contributed to a better understanding of the higher-level processing of visualizations, including comprehension [bateman2010useful], engagement [borgo2012empirical], memorability [borkin2015beyond], and biases [dimara2018task]. Many of these empirical findings have been distilled into design guidelines that are now widely referenced in practitioner resources [choi2023vislab]; examples include using bar charts instead of pie charts for comparing values or including a zero baseline. While empirical research continues to expand and refine such guidelines, it remains unclear how well research-driven guidance is reflected in the guidelines practitioners actually use.
In this study, we investigate the research-practice gap in visualization design guidelines, aiming to understand their interplay better and promote evidence-based design practices. As the field of visualization increasingly influences decision-making across society, it is crucial to bridge this gap to help practitioners make informed choices for creating truthful and functional visualizations. Our approach mirrors initiatives in fields such as medicine, education, and policy-making, where the shift has been from relying on tradition, intuition, and anecdotal experience to making decisions based on solid research findings [evidpractice, bogenschneider2011evidence, davies1999evidence, sackett1996evidence]. This movement also relates to intermediary efforts, often referred to as translational science [woolf2008meaning], which focus on converting research results in research into practical implementations [norman2010research]. Although translation can go in both directions, our focus is specifically on examining the design guidelines accessible to practitioners and assessing how well they align with empirical research findings.
Our specific research questions include:
@@
  • How do practitioners and researchers perceive the research-practice gap in visualization design guidelines?
@@
We employed a two-phase mixed-methods approach to explore these questions. We first collected data visualization design guidelines in the wild by conducting a keyword-based web search, and then conducted a qualitative content analysis to characterize their attributes. Our findings revealed a concentration of guidelines on basic charts and task types and identified three distinctive categories of the guidelines: problem-solving, enhancement, and alternative suggestion. To understand the agreement and discrepancies between guidelines and empirical research, we analyzed how they are aligned with each other; we focused on empirical studies that evaluate user performance across various visualization designs [lam2011empirical]. The alignment analysis revealed inconsistencies between mapped guidelines and their associated empirical evidence, highlighting a disconnect between research findings and practitioner-facing guidelines. Unmapped studies suggest areas for new guidelines, while unsupported guidelines offer opportunities for empirical research.
@@
The content analysis of the guideline landscape provided insights into the available guideline knowledge in practice, but it did not capture the lived experiences and perspectives of researchers and practitioners regarding the knowledge gaps. To gain a more comprehensive understanding, we supplemented this with surveys and interviews. Practitioners were mainly recruited from the Data Visualization Society [datavissociety], while researchers were recruited through community mailing lists. A total of 33 practitioners and 36 researchers filled out our survey questionnaire on guidelines, empirical studies, and the challenges of transferring knowledge between the two communities. The survey results revealed similarities and differences in their access to guidelines, perceptions of their usefulness, and the challenges they encountered, while also indicating a shared recognition of the gap in transferring research into design guidelines [parsons2021understanding]. In follow-up interviews with ten selected participants from each community, we gained deeper insights into their experiences and struggles with guidelines, as well as their attitudes and perceived causes of the knowledge gap and ideas for addressing it.
@@
Our primary contribution lies in enhancing the understanding of the design guideline knowledge gap between research and practice, paving the way for evidence-informed design decision-making. We synthesized our findings into a range of strategies and considerations, detailed in the discussion section, aimed at bridging the guideline-focused knowledge divide. These strategies include establishing consistent vocabularies for data visualizations, creating a triadic model that integrates guidelines, empirical evidence, and contextual examples, and improving guideline accessibility through detailed metadata annotation and template development. We also highlight the importance of fostering collective efforts in knowledge brokerage [ward2009knowledge], essential for the growth and evolution of the field.
@@ II Background: Research, Practice, & Education
The field of data visualization has always been closely tied to practical needs and applications. Widely used software tools and programming toolkits, such as Tableau [tableau, stolte2002polaris], D3 [d3], and Vega-Lite [vegalite], began as research projects, as did visualization techniques such as treemaps [johnson1998tree] and parallel coordinates [inselberg1990parallel]. Researchers have also devoted much effort to understanding the challenges and struggles faced by practitioners in deriving new research insights [joyner2022visualization, bigelow2014reflections] and building new visualization tools for practitioners [kim2016data, liu2018data]. Furthermore, the growing sub-fields of visual analytics [cook2005illuminating] and design studies [sedlmair2012design] directly address problems faced by domain experts across disciplines.
@@
Similarly, the visualization research community has developed theories and frameworks over the past few decades that provide established vocabularies for reading and creating visualizations, such as visual encoding channels, data types, and analytical tasks [brehmer2013multi, munzner2014visualization]. To evaluate the usefulness of a visualization design, researchers conduct experiments that compare design alternatives. The results of these experiments provide guidance for selecting effective visual encoding channels [cleveland1984graphical, heer2010crowdsourcing] and determining their effectiveness for different tasks [kim2018assessing, saket2018task]. Some of these results are integrated into existing tools that automate visual encoding design [tableau, voyager, hu2019vizml]. Although these automated tools suffice for exploratory analysis, designers must still consider a range of design factors to ensure the successful and engaging communication of data, such as selecting appropriate color schemes and determining axis ranges.
@@
Researchers have developed pedagogical approaches to teaching data visualizations, such as the Five Design-Sheet Methodology [roberts2015sketching] and VisItCards [he2016v]. Other studies have examined data visualization literacy, which is defined as the ability to read information encoded in data visualizations and create them to answer questions about the data [lee2016vlat, borner2019data, alper2017visualization, boy2014principled, ge2023calvi]. However, these works in pedagogy and literacy tend to prioritize higher-level design processes and chart-reading tasks, often placing less emphasis on practical knowledge like actionable design guidelines.
@@
Several recent studies have attempted to address more practical design needs. Diehl et al. launched VisGuides, a forum that provides practitioners with a central place to ask questions about the validity and quality of their visualizations and discuss guidelines around visualization design [diehl2018visguides]. Wang et al. proposed a cheatsheet structure to introduce a visualization technique, including its construction process, similar techniques, and potential issues [wang2020cheatsheet]. On the other hand, researchers have discussed theoretical models for connecting with empirical studies [brown2018tagguideline] with guidelines and using guidelines to describe relationships between various stages of the design process [meyer2014nestedblock].
@@
Similar efforts have also been made by practitioners as well. Resources such as From Data to Viz and DataVizProject provide user interfaces to select charts based on different data types and tasks [datatoviz, datavizproject]. Companies publish design systems that often include recommendations on visualization design [adobedesignsystem, googledesignsystem], although they mostly focus on styling guidelines to match their brands.
@@
Currently, visualization design guidelines are dispersed across different resources, often through personal blogs and published books. It remains unclear how, or even if, practitioners learn critical design knowledge from these guidelines. Furthermore, researchers often convey research outcomes through blogs [multipleviews], podcasts [datastories], or social media in an ad-hoc manner. While recent studies have begun investigating the gap between research and practice, they mostly focus on understanding holistic design processes rather than acquiring fundamental data visualization design knowledge [parsons2021understanding, zhang2022visualization].
This paper aims to understand the design guideline knowledge gap between visualization design practice and research. Bridging this gap is essential because it enables practitioners to leverage the latest research to make more informed decisions, resulting in better data visualizations. In addition to fields including medicine and biology, people in a related field of human-computer interaction, have recognized this issue as well and studied how to better apply research findings to practice, a process known as translational research [woolf2008meaning].
For instance, researchers conducted interviews with practitioners to understand their barriers to research findings [colusso2017translational] and to propose new models to tackle the research-practice gap [colusso2019translational, beck2018theory], mostly focusing on the domain of user experience design. Other researchers have examined how specific research frameworks are applied by user experience designers [velt2020translations, remy2015bridging]. Other studies have focused on specific subdomains such as AI in UX [yildirim2023investigating], human-robot interaction [cila2024bridging] interaction design for children [kawas2021translational].
We focus on data visualization which is uniquely different from these fields. The design of effective visualizations can be grounded in perceptual and cognitive studies through systematic investigations based on a well-defined design space, including encoding, data types, and tasks, offering a distinct perspective on alleviating the gap between research and practice. In this work, we employ a range of mixed methods, including content analysis, surveys, and interviews, engaging both practitioners and researchers, to investigate the knowledge gap in guidelines.
@@ III Guideline and Empirical Study Alignment
To better understand the research-practice gap in visualization design guidelines, we first examined how existing practitioner-facing guidelines align with empirical research findings. This analysis helps reveal whether current guidelines are well-supported by research evidence, identify areas of mismatch, and uncover opportunities for new guideline development.
@@ III-A1 Collecting guidelines
To ensure comprehensiveness in our collection, we extensively consulted a variety of guideline resources in practical settings, such as industry resources and online platforms. We first used web search by using a combination of different keywords, such as data visualization, guidelines, and similar synonyms, and went through up to five pages of the top results. We also looked at known resources such as blogs (e.g., DataWrapper [datawrapper]) and websites (e.g., from Data to Viz [datatoviz], Data Journalism [datajournalism]). Finally, we consulted the guideline collection curated by the Data Visualization Society [datavissociety], which provided additional blogs, websites, and books (e.g., How Charts Lie by Cairo [cairo2019charts]). Our search concluded once we reached a saturation of overlapping and duplicate guidelines, totaling 1,279 guidelines.
Next, we applied inclusion and exclusion criteria to serve our purpose of analyzing visualization guidelines in practice. For instance, we only considered guidelines relevant to data visualization design, adopting the definition from Chen et al. [chen2017pathways]; i.e., a guideline describes a process or a set of actions that may lead to a desired outcome or, alternatively, actions to be avoided to prevent an undesired outcome. Consequently, we excluded guidelines offering narrow styling advice, such as increasing font size. We also disregarded broader recommendations, such as making visualizations interactive. Similarly, we excluded generic chart suggestions, such as using a bar chart for a single series data, that are too evident to be considered guidelines. We also consolidated guidelines that significantly overlapped in content or purpose. After the filtering and aggregation process, we obtained the final collection of 390 guidelines from 32 diverse sources. Of these, 49% were derived from blog articles, 39% from style guidelines, and the remaining 12% from books.
@@
We focused on practitioner-facing sources to capture the guidelines actually consumed by the target audience, distinguishing them from the academic source material. Findings from Section IV), along with prior literature [colusso2017translational], confirm that practitioners rarely consult academic research papers.
Why we did aggregated similar guidelines
We aimed to avoid overestimating the results of our alignment analysis, which examines the extent to which guidelines align or mis-align with empirical research by giving equal weight to each guideline. However, analyzing the overlap in the frequency of each guideline could reveal which ones are most commonly used in practice, though this is not the primary focus of this paper.
Why we only considered certain types of guidelines
Design knowledge is a broad concept that encompasses anything related to design. To narrow our scope, we focused on guidelines using the definition provided by Chen et al. However, guidelines themselves can also be broadly defined, so we refined our focus to those that contain meaningful design knowledge that is specific to data visualization. For instance, a chart type predetermines the kinds of data that can be represented, and thus it may not require empirical support. On the other hand, while guidelines can be based on experience, they may still warrant empirically-oriented explanations (e.g., making a title engaging  [borkin2015beyond]).
@@ III-A2 Collecting empirical studies
We used a combination of automated and manual searches to collect empirical papers. First, we initially used a keyword search (e.g., visualization, evaluation, study) on Scopus to collect papers from ACM CHI and EuroVis over the past decade. We collected the same data for IEEE VIS from VisPub [isenberg2016vispubdata]. Expanding on this initial collection, we supplemented it with recent papers published in ACM CHI, IEEE VIS, and Eurovis from 2021 to 2022 to ensure the dataset reflects the latest research. We also manually included relevant past papers that were not captured by our initial search from additional sources, including VisPerception [visperception] and a recent survey of perception-based visualization studies [quadri2021survey]; this process aided in gathering influential works published prior to 2011, such as the notable graphical perception experiment conducted by Cleveland & McGill [cleveland1984graphical]. Finally, two researchers cleaned up the data collection by removing any papers that did not present original empirical research and by only including studies that evaluate user performance on different visualization designs based on the characterization in a prior work [lam2011empirical] i.e., we excluded algorithm performance and system usability studies. The final collection of empirical studies consists of 235 research papers, spanning from 1911 to 2022. A detailed breakdown of the papers by venues and years is available in the supplementary material.
@@ III-B Analyzing Process
To facilitate content analysis [potter1999rethinking] of the guidelines, two junior researchers compiled short descriptions from guideline articles into a spreadsheet. A lead researcher initially reviewed these descriptions and, when necessary, consulted the source articles. This process was to establish initial codes and code values, including chart types, data types, task types, and chart elements. The two junior researchers then independently coded the data using the initial codebook. They expanded it as necessary, adding new codes related to purpose, problem, approach, and specific features like color, animation, 3D, and interaction, consulting with the lead researcher to resolve any questions or conflicts in coding. Lastly, the lead researcher thoroughly reviewed the final codes.
@@
We derived the one-sentence purpose of each study out of its abstract using GPT4 to facilitate the following alignment process. We initially attempted to codify aspects such as chart types and experiment methods. However, given the comprehensive empirical study landscape already explored in recent studies [kosara2016empire, zeng2023review, quadri2021survey], we ultimately decided not to perform a detailed codification of the empirical studies. This approach allowed us to maintain a high-level analysis that aligns with our study’s strategic objectives, ensuring a focused and manageable research endeavor. Nonetheless, exploring finer-grained, code-level alignment in future work beyond our current entity-level alignment would be beneficial.